OpenStudy (anonymous):

If 17y-24y+13y=210+y, what is the value of y?

OpenStudy (texaschic101):

17y - 24y + 13y = 210 + y --- combine like terms on the left side 6y = 210 + y -- subtract y from both sides 6y - y = 210 5y = 210 y = 210/5 y = 42 you can even check your answer 6y = 210 + y 6(42) = 210 + 42 252 = 252 (correct)

OpenStudy (anonymous):

it says it's incorrect

OpenStudy (texaschic101):

Is the problem written correctly ?

OpenStudy (anonymous):

I checked it and it is written correctly

OpenStudy (texaschic101):

I am sorry but I do not see a mistake in my calculations

OpenStudy (anonymous):

ok well that's ok, can anybody else help me